Bonita invested R40 000 in order to have R56 000 available in 30 months' time. The yearly rate, compounded semi-annually, is

[1] 7,21%.

[2] 16,00%.

[3] 8,00%.

[4] 14,41%.

[5] 13,92%.


1
Expert's answer
2021-06-14T19:15:50-0400

GghhhhhInvestment = 40000

Amount in 30 months = 56000

Months = 30

Hal yearly terms (n) = 30/6 =5

Interest rate for 6 months = r


Therefore;

"=> 40000\times(1+r)^n = 56000"


"=> (1+r) ^5 = \frac{56000}{40000}"


"=> 1+r = 1.40^{(\frac{1}{5})}"


"=> r = 1.0696 - 1"


"=> r = 0.0696"


"=> r = 6.96\%"


Therefore, Interest rate per annum


"6.96\%\times2 =13.92\%"